saniyawrenn:

Which equation could be used to find the number of minutes, m, in h hours? M= 60h M= 60/h M= 60+h M= h/60

jhonyy9:

so i think everybody know that 1 hour has 60 minutes so in this way @TWALKER16 how you thought this above wrote pls. ? so from what you wrote above result that 1 minute equal 60 hours this is true ?

saniyawrenn:

@jhonyy9 I think that they meant it this way. If you think about it, H=Hours. There is 60 minutes in an hour. You multiply. I think idk im stuck

jhonyy9:

ok look please step by step M= 60h M= 60/h M= 60+h M= h/60 the first one ?

jhonyy9:

so pls. collaborate